how many grams of salt must be added to 40grams of a 28% salt solution to make it 40%?
Answer by robertb(5830)

Let x = # grams salt that must be added. Then
0.28*40 + x = 0.40(40+x),
11.2 + x = 16 + 0.40x,
0.60x = 4.8,
x = 8 grams of salt, must be added.